OpenStudy (anonymous):

how do i enter int (1,sqrt x) of (2t-1)/(t+2)dt

OpenStudy (anonymous):

it would be\[\int\limits_{1}^{\sqrt{x}}\frac{2t-1}{t+2}dt\] In the equation editor it would look like: int_{1}^{sqrt{x}}frac{2t-1}{t+2}dt Pretty complicated, right? :)